Accelerator Update for July 25, 2008 - July 28, 2008
Friday July 25
The day shift began with Operations monitoring store 6317 (93.30E30) and the Recycler (RR) stash (122.5E10), with the Antiproton Source stacking (28.6E10), with the Main Injector, Booster, and Linac operating normally, with the SciBooNE / MiniBooNE (BNB), NuMI, and Meson MTest experiments taking beam.
MTest turned off at 12:10 PM.
At 2:06 PM a water leak was found on Booster GMPS #2 power supply, beam was halted and investigation/repair began.
Beam returned at 3:23 PM.
Operations established store 6318 at 10:24 PM with an initial luminosity of 290.76E30.
Beam returned to MTest at 10:29 PM.
D0 began experiencing network problems at 11:20 PM, they are unable to take data.
Saturday July 26
The midnight shift began with operations monitoring store 6318 (216.57E30) and the stash (81.5E10), with Pbar stacking (27.7E10), with the BNB, NuMI, and MTest experiments taking beam.
D0 began taking data once again at 6:05 AM. The problem was traced to a small power outage (caused by a blown fuse) to their level 3 computer cluster. FESS electrician was to replace the fuse.
MTest turned off at 12:21 PM.
Store 6318 was terminated at 3:45 PM.
Operations established store 6319 at 5:56 PM with an initial luminosity of 294.26E30.
Beam returned to MTest at 8:00 PM.
Sunday July 27
The midnight shift began with operations monitoring store 6319 and the stash, with Pbar stacking, with the BNB, NuMI, and MTest experiments taking beam.
At 5:07 AM the E-Cool frontend stopped responding to a node poll. Two software reboots were tried, both were unsuccessful. Pbar transfers were put on hold and experts were called to investigate.
MTest turned off at 10:04 AM.
At 12:15 PM the E-Cool frontend problems were repaired and the Recycler was recovered from various problems associated with accessing MI-31 to facilitate the frontend repair.
Store 6319 was terminated at 2:02 PM.
Operations put the Tevatron (TeV) into shot setup at 2:12 PM.
Operations established store 6320 at 4:19 PM with an initial luminosity of 296.63E30.
Beam returned to MTest at 8:21 PM.
Monday July 28
The midnight shift began with operations monitoring store 6320 and the stash, with Pbar stacking, with the BNB, NuMI, and MTest experiments taking beam.
At 4:03 AM, Booster BRF2 and BRF3 tripped off due to an LCW leak. Booster was turned off for an access to investigate.
Booster beam returned at 5:29 AM. A water leak was found on BRF2 and water was valved off to the cavity. Booster is currently running without BRF2.
At 8:26 AM, CDF reported a high voltage crate problem which is causing them to mark their data as bad.
The Plan for Monday
The plan for today is to terminate the store at 9:15 AM to let CDF repair their high voltage crate problem. Later today put in a new Tev store, and resume normal stacking, stashing, and storing.
Comments and Suggestions
|last modified 07/28/2008 email Fermilab|
|Security, Privacy, Legal|